A weather balloon is inflated to a volume of 25.6 L at a pressure of 735 mmHg and a temperature of 27.4 ∘C∘C. The balloon rises in the atmosphere to an altitude where the pressure is 375 mmHg and the temperature is -15.9 ∘C. Assuming the balloon can freely expand, calculate the volume of the balloon at this altitude.


V balloon = --- L





1
Expert's answer
2021-05-31T02:07:20-0400

(P1 × V1) / T1 = (P2 × V2) / T2


therefore: "V2 =\frac {(P1 × V1 × T2)} {(T1 × P2)}"


P1 = 732 mmHg

V1 = 27.9 L

T2 = -13.6C = 259.55K

T1 = 30.1C = 303.25K

P2 = 385 mmHg


V2 = (732 × 27.9 × 259.55) / (303.25 × 385)

V balloon= 45.4 L